Home‎ > ‎Crew 161‎ > ‎

0113-DailySummaryReport

MDRS Daily Summary Report
Sol 4 1/13/2016

North Pinto Hills and Radio Problems.

Executive Officer Reporting

Mission Status: Research projects are on track and mission is going on smoothly. Glitches with communication gadgets

Sol Activity Summary: The highlight of SOL-4 was EVA #3 which was led by Commander Divya. EVA #3 explored the areas around the North Pinto Hills. Commander Divya collected samples from three depths, measured pH values, moisture content and the shear capacitance at the different depths around the North Pinto Hills. Crew Astronomer, captain Marmot was on ground to take beautiful shots of this EVA.

Look Ahead Plan: An EVA is planned for tomorrow to test radio communications range and take some photos.

Anomalies in work: Radios seem to have a very short range. Less than 3 miles away from the HAB.

Weather: The Martian sky was bright and clear today.

Crew Physical Status: All crew members are hale and hearty!

EVA: Three soil samples were collected from the North Pinto Hills Mars during today’s EVA. EVA duration exceeded allocated time. Radio communication with HABCOMM was lost within 4 miles of HAB, however EVA was not aborted. Commander Divya will be studying the chemical composition of collected samples.

Reports to be filed:
Commanders Report
EVA #3 report
Daily Summary Report
Astronomy report
Journalist Report
Biologist Report
HSO report
Engineering Report

Support Requested: Gasoline refill requested. Please advise on Radio communications range.